John Brown, career counselor, needs a program that will sort student names. Write a program that will sort a list of 10 student names that is inputted by the user. Save the program as studentsort.py. Comment the program as to which sort method was used.
Answer #1

I have used insertion sort ..

Here is code:

def sortList(list):

for i in range(1, len(list)):

key = list[i]

j = i - 1

while j >= 0 and key < list[j]:

list[j+1] = list[j]

j -= 1

list[j+1] = key

return list



l = ["Otto Denekamp","Mac Andrin","Shandie Fuggles","Jenna Orknay","Flore Ezzell","Si Fernao","Ediva Edgley","Sayers Albisser","Trenna Wabersich","Morgana Symons"]

print("Array list before sort :")

print(l)

print(" Array list after sort :")

print(sortList(l))


Output: